# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ Abstract
In recent years, multi-agent teaming has garnered considerable interest since complex objectives, such as intelligence, surveillance, and reconnaissance, can be divided into multiple cluster-level sub-tasks and assigned to a cluster of agents with the appropriate functionality. Yet, coordination and information dissemination between clusters may be necessary to accomplish a desired objective. Distributed consensus protocols provide a mechanism for spreading information within clustered networks, allowing agents and clusters to make decisions without requiring direct access to the state of the ensemble. Hence, we propose a strategy for achieving system-wide consensus in the states of identical linear time-invariant systems coupled by an undirected graph whose directed sub-graphs are available only at sporadic times. Within this work, the agents of the network are organized into pairwise disjoint clusters, which induce sub-graphs of the undirected parent graph. Some cluster sub-graph pairs are linked by an inter-cluster sub-graph, where the union of all cluster and inter-cluster sub-graphs yields the undirected parent graph. Each agent utilizes a distributed consensus protocol with components that are updated intermittently and asynchronously with respect to other agents. The closed-loop ensemble dynamics is modeled as a hybrid system, and a Lyapunov-based stability analysis yields sufficient conditions for rendering the agreement subspace (consensus set) globally exponentially stable. Furthermore, an input-to-state stability argument demonstrates the consensus set is robust to a class of perturbations. A numerical simulation considering both nominal and perturbed scenarios is provided for validation purposes.
I Introduction
I-A Motivation
@@
The Internet of Things (IoT) and advancements in data handling within network systems have created a paradigm shift in how the internet and distributed platforms are used today [1]. Consumer electronics and household devices such as cameras, watches, televisions, and refrigerators are now equipped with WiFi or LAN functionality granting the user access to nearly instantaneous data and wider control of their environment [2]. Connected devices, like those listed above, possess a natural graphical structure called a cluster-based network [3] or small-world network [4, 5, 6]. Although sophisticated, the technology enabling current small-world networks will need to evolve so that it can address imminent challenges. For instance, as more devices gain internet connectivity and transmit more data due to improved capability, the resulting small-world networks will increase in complexity and require enhanced data handling and control protocols to meet demands. Bottlenecks introduced by finite network resources, e.g., bandwidth and throughput, and the need to perform control using limited information are ever-present constraints for multi-agent systems (MASs), especially those consisting of platforms with dynamics, such as vehicles.
@@
Given their wide applicability and utility in coordination [7], we are interested in distributed consensus protocols for MASs. Specifically, we focus on consensus over clustered networks while utilizing intermittent and asynchronous output feedback. Data-handling protocols that embrace intermittency and asynchrony can better accommodate large-scale networks and data-dense exchanges than those using continuity and synchrony. Moreover, output feedback is an example of limited information. In practice, certain components of cyber-physical systems may operate at a constant rate, for example, electromechanical sensors and digital processors. Consequently, some information needed for control may be available periodically, motivating the development of controllers that leverage intermittent feedback to seamlessly integrate with such components. Moreover, intermittency can promote efficient resource utilization relative to legacy systems requiring continuous or periodic updates. The agents should also accommodate asynchronous information since delays in data propagation across networks and different operating tempos between agents occur naturally in distributed systems.
I-B Background
@@
The consensus problem for MASs, in both the continuous-time and discrete-time settings, has been well-studied in [8]. Furthermore, the survey papers in [9, 10, 11] provide additional information on MASs and consensus. Recently, however, there has been significant research on consensus where the communication between agents and control updates within agents are periodic, intermittent, and/or event-triggered [12, 13, 14, 15, 16, 17, 18, 19]. These works were motivated by the efficient use of limited resources and the implementation challenges inherent to cyber-physical systems caused by, for example, sensor sampling rates, radio broadcast rates, and actuator constraints. Yet, there is a lack of research focusing on consensus over clustered networks, which is somewhat surprising since it is natural to solve complicated problems with cooperative teams of specialized agents. In fact, clustering in network systems is quite common in engineering applications [20, 21, 22, 3, 23, 24].
@@
The results in [25, 26, 27, 28, 29, 30, 31] delve into the control of MASs over clustered networks. In [25], the authors develop a distributed controller enabling $p$-cluster consensus for a MAS of identical linear time-invariant (LTI) systems coordinating over a static or switching graph. Given a graph encoding a MAS whose vertex set is partitioned into $p$ elements (i.e., clusters), $p$-cluster consensus is said to occur whenever the agents in each cluster achieve consensus, and the consensus value between clusters is distinct. Note that $p$-cluster consensus is accomplished using continuous state feedback. In [26], the authors investigate the formation control problem for clustered networks of identical LTI systems. The authors develop a distributed controller that enables $M$ clusters to assemble distinct rigid formations using intermittent state feedback while compensating for communication delays.
@@
The result in [27] presents a distributed consensus controller for a clustered network of identical LTI systems, where each agent leverages a mixture of continuous and intermittent output feedback. Each cluster is assigned a single leader, and, within each cluster, all agents may continuously communicate output information with their neighbors. However, only the leaders of each cluster may communicate intermittently with neighboring leaders, and this is the means by which information is spread across clusters. In addition, all leaders communicate with their neighboring leaders at the same isolated times. A Luenberger-like observer produces state estimates from output information, which are then used for control. Albeit not written, the analysis in [27] renders the agreement subspace for the entire ensemble globally exponentially stable. Leveraging analogous arguments and cluster structures found in [27], the result in [28] develops a distributed controller enabling a clustered network of identical LTI systems to assemble a single rigid formation using a mixture of continuous and intermittent state feedback. In [29], the authors build on the result of [25] and identify the general graph property that supports $p$-cluster consensus independent of coupling strength.
I-C Comparison to Previous Work
@@
In [30], we build on [27] and develop a distributed controller for a clustered network of identical LTI systems that enables consensus across the entire ensemble. Like [27], communication within clusters is continuous, and communication between clusters is intermittent. Nevertheless, unlike [27], there is no leader-follower structure in [30], and more general partitions of the communication graph are considered. As a result, more than one agent in a cluster can have a neighbor in a different cluster. Moreover, in [30], two pairs of distinct clusters may exchange information at different rates (i.e., asynchronously), unlike [27]. Still, it is important to observe that [27] considers output feedback while [30] only considers state feedback.
@@
Our result in [31] is an extension of [30]. More precisely, [31] develops a distributed controller for a clustered network of identical LTI systems that enables ensemble-wide consensus using output feedback. Similar to the results in [27] and [30], communication in clusters is continuous, while communication between clusters is intermittent. Yet, rather than broadcast state information between neighbors as in [30], state estimates are broadcast instead, which are obtained using a Luenberger-like observer. This observer intermittently samples the output of an agent and leverages model knowledge and the control input to compute a state estimate.
@@
Unfortunately, the stability analysis in [31] suffers from an artificial defect. To demonstrate the desired set $\mathcal{A}$—which has the agreement subspace as a component—is globally exponentially stable, the stability analysis required all Luenberger-like observers to undergo synchronous jumps. The price to pay for this mathematical convenience is that all agents must synchronously sample their output, which is unreasonable for large-scale networks.
I-D Contributions
Motivated by our previous work in [30, 31], we introduce a distributed consensus controller for a clustered network of identical LTI systems that uses intermittent and asynchronous output feedback. A connected, undirected, and static communication graph that is decomposed into connected, undirected, and static sub-graphs, with union returning the parent graph, defines the notion of clustered network within this work. These sub-graphs are separated into two groups, that is, clusters and inter-clusters, which are precisely defined in Section II. Each agent and inter-cluster is assigned a timer that dictates when measurement and control update events take place.
Within each cluster, each agent intermittently measures the outputs of their same-cluster neighbors and themselves when instructed by their timer mechanism. Similarly, within an inter-cluster, each agent intermittently measures the outputs of their same-inter-cluster neighbors and themselves when instructed by the inter-cluster timer mechanism. Such measurements may occur asynchronously between agents and inter-clusters. Thus, the measurement of output information is completely intermittent throughout the entire network, unlike [27, 30] and [31] which have components that use continuous communication to obtain output information.
The outputs are used to impulsively reset estimators, which evolve in an open-loop fashion during flows. The estimators replace the Luenberger-like observers from [31] and the need for explicit model knowledge in the controller. In addition, the controller utilizes the estimator outputs to facilitate consensus in the states of the LTI systems. To capture the continuous-time and discrete-time behaviors of the closed-loop ensemble system, we construct a hybrid model of the form in [32]. We also transform the consensus objective into a set stabilization problem for hybrid systems. A set $\mathcal{A}$, encoding the consensus subspace, is demonstrated to be globally exponentially stable using a Lyapunov-based stability analysis. The nominal well-posedness of the proposed hybrid system implies that the set $\mathcal{A}$ is robust to vanishing perturbations. An input-to-state stability argument for hybrid systems is also provided to show that the set $\mathcal{A}$ is robust to perturbations that are Lebesgue measurable and locally essentially bounded for the closed-loop hybrid system. Such an input-to-state stability argument cannot be found in [27, 30] and [31]. Lastly, we would like to stress that this work, as well as our previous works [30, 31], provide an alternative method of defining clustered networks and their associated sub-graphs when compared to the works in [25, 26, 27, 28, 29]. This particular construction is generic and facilitates the use of standard analysis techniques intended for traditional MASs. The theoretical development is validated in simulation, where we provide an online path planner for a rendezvous application.
I-E Organization
@@
Section II provides information on notation, graphs, cluster and inter-cluster sub-graphs, and hybrid systems. Section III describes the specific problem we are aiming to solve. The proposed solution and closed-loop hybrid inclusion are presented in Section IV. In Section V, we provide the main results of the paper in the form of sufficient conditions for the nominal system and, in Section VI, we extend these results to consider the case of perturbations. Lastly, we illustrate the results through numerical simulations in Section VII.
II Preliminaries
II-A Notation
@@
Let $\mathbb{R}$ and $\mathbb{Z}$ denote the set of reals and integers, respectively. For a constant $a\in\mathbb{R}$, let $\mathbb{R}_{\geq a}\triangleq[a,\infty)$, $\mathbb{R}_{>a}\triangleq(a,\infty)$, $\mathbb{Z}_{\geq a}\triangleq\mathbb{R}_{\geq a}\cap\mathbb{Z}$, and $\mathbb{Z}_{>a}\triangleq\mathbb{R}_{>a}\cap\mathbb{Z}$. For $p,q\in\mathbb{Z}_{>0}$, the $p\times q$ zero matrix and the $p\times 1$ zero column vector are respectively denoted by $0_{p\times q}$ and $0_{p}$. The $p\times p$ identity matrix and the $p\times 1$ column vector with all entries being one are denoted by $I_{p}$ and $1_{p}$, respectively. The Euclidean norm of $r\in\mathbb{R}^{p}$ is denoted by $\|r\|\triangleq\sqrt{r^{\top}r}$. For $x,y\in\mathbb{R}^{p}$, the inner product between $x$ and $y$ is denoted by $\langle x,y\rangle\triangleq x^{\top}y$. For $M\in\mathbb{Z}_{\geq 2}$, let $[M]\triangleq\{1,2,...,M\}$. The Kronecker product between $A\in\mathbb{R}^{p\times q}$ and $B\in\mathbb{R}^{u\times v}$, where $u,v\in\mathbb{Z}_{>0}$, is denoted by $A\otimes B\in\mathbb{R}^{pu\times qv}$. The maximum and minimum eigenvalues of a real symmetric matrix $A\in\mathbb{R}^{n\times n}$ are denoted by $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$, respectively. The block diagonal matrix with general blocks $G_{1},G_{2},...,G_{p}$ is denoted by $\text{diag}(G_{1},G_{2},...,G_{p})$. The distance of a point $r\in\mathbb{R}^{p}$ to the set $S\subset\mathbb{R}^{p}$ is given by $|r|_{S}\triangleq\inf\{\|r-s\|\colon s\in S\}\in\mathbb{R}_{\geq 0}$; furthermore, define the set addition property as $r+S\triangleq\{r+s\in\mathbb{R}^{p}\colon s\in S\}$. For a collection of vectors $\{z_{1},z_{2},...,z_{p}\}\subset\mathbb{R}^{q}$, let $(z_{k})_{k\in[p]}\triangleq[z_{1}^{\top},z_{2}^{\top},...,z_{p}^{\top}]^{\top}\in\mathbb{R}^{pq}$. Similarly, for $x\in\mathbb{R}^{p}$ and $y\in\mathbb{R}^{q}$, let $(x,y)\triangleq[x^{\top},y^{\top}]^{\top}\in\mathbb{R}^{p+q}$. For any sets $A$ and $B$, a function $f$ of $A$ with values in $B$ is denoted by $f\colon A\to B$, whereas $f\colon A\rightrightarrows B$ refers to a set-valued function $f\colon A\to 2^{B}$. The $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N}$, for $N\in\mathbb{Z}_{>0}$, is denoted by $\mathtt{e}_{k}$. Similarly, let $\tilde{\mathtt{e}}_{k}$ denote the $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N-1}$. The range (i.e., image) of a matrix $A\in\mathbb{R}^{N\times N}$ is denoted by $\text{Range}(A)\triangleq\{y\in\mathbb{R}^{N}\colon y=Ax\text{ for some }x\in\mathbb{R}^{N}\}$. For any $x\in\mathbb{R}$, let $\lfloor x\rfloor\triangleq\sup\{c\in\mathbb{Z}\colon c\leq x\}$.
II-B Graphs

### v3

# 2]organization=Air Force Research Laboratory, Space Vehicles Directorate, city=Kirtland AFB, state=NM, postcode=87117, country=USA

@@ Abstract
Distributed consensus protocols provide a mechanism for spreading information within clustered networks, allowing agents and clusters to make decisions without requiring direct access to the state of the ensemble. In this work, we propose a strategy for achieving system-wide consensus in the states of identical linear time-invariant systems coupled by an undirected graph whose directed sub-graphs are available only at sporadic times. Within this work, the agents of the network are organized into pairwise disjoint clusters, which induce sub-graphs of the undirected parent graph. Some cluster sub-graph pairs are linked by an inter-cluster sub-graph, where the union of all cluster and inter-cluster sub-graphs yields the undirected parent graph. Each agent utilizes a distributed consensus protocol with components that are updated intermittently and asynchronously with respect to other agents and inter-clusters. The closed-loop ensemble dynamics is modeled as a hybrid system, and a Lyapunov-based stability analysis yields sufficient conditions for rendering the agreement subspace (consensus set) globally exponentially stable. Furthermore, an input-to-state stability argument demonstrates the consensus set is robust to a large class of perturbations. A numerical simulation considering both nominal and perturbed scenarios is provided for validation purposes.
keywords:
[…]
1 Introduction
1.1 Motivation
@@
The Internet of Things (IoT) and advancements in data handling within network systems have created a paradigm shift in how the internet and distributed platforms are used today [Li.Hu.ea2021]. Consumer electronics and household devices such as cameras, watches, televisions, and refrigerators are now equipped with WiFi or LAN functionality, granting the user access to nearly instantaneous data and wider control of their environment [Chang.Srirama.ea2019]. Connected devices, like those listed above, possess a natural graphical structure called a cluster-based network [Wang.Xiao2006] or small-world network [Zaidi2013, Telesford.Joyce.ea2011, Watts.Strogatz1998]. Although sophisticated, the technology enabling current small-world networks will need to evolve so that it can address imminent challenges. For instance, as more devices gain internet connectivity and transmit more data due to improved capability, the resulting small-world networks will increase in complexity and require enhanced data handling and control protocols to meet demands. Bottlenecks introduced by finite network resources, e.g., bandwidth and throughput, and the need to perform control using limited information are ever-present constraints for multi-agent systems (MASs), especially those possessing dynamics, such as vehicles.
@@
Given their wide applicability and utility in coordination [Mesbahi.Egerstedt2010], we are interested in distributed consensus protocols for MASs. Specifically, we focus on consensus over clustered networks while utilizing intermittent and asynchronous output feedback. Data-handling protocols that embrace intermittency, asynchrony, and limited information are better suited for large-scale networks than policies requiring periodically and synchronously derived full state feedback. In practice, certain components of cyber-physical systems may operate at a constant rate, for example, electromechanical sensors and digital processors. Hence, some information needed for control may be available periodically, motivating the development of controllers that can seamlessly integrate with such components. Intermittency can also promote the efficient use of limited computation and communication resources.
1.2 Background
@@
The consensus problem for MASs, in both the continuous-time and discrete-time settings, has been well-studied in [OlfatiSaber.Fax.ea2007]. Furthermore, the survey papers in [Qin.Ma.ea2017a, Li.Tan2019, Dorri.Kanhere.ea2018] provide additional information on MASs and consensus. Recently, however, there has been significant research on consensus where the communication between agents and control updates within agents are periodic, intermittent, and/or event-triggered [Ding.Han.ea2017, Nowzari.Garcia.ea2019, Ge.Han2016, Phillips.Sanfelice2024, Zegers.Guralnik.ea2022a, Phillips.Sanfelice2019a, Zegers.Guralnik.ea2024, Fan.Yang.ea2016]. These works were motivated by the efficient use of limited resources and the implementation challenges inherent to cyber-physical systems caused by, for example, sensor sampling rates, radio broadcast rates, and actuator constraints. Yet, there is a lack of research focusing on consensus over clustered networks, which is somewhat surprising since it is natural to solve complicated problems with cooperative teams of specialized agents. In fact, clustering in network systems is quite common in engineering applications [Wirasanti.Ortjohann.ea2012, Cheng.Scherpen2018, SanchezGarcia.Fennelly.ea2014, Wang.Xiao2006, Bandyopadhyay.Coyle2004, Jayawardene.Herath.ea2020].
@@
The results in [Qin.Yu2013, Ge.Han.ea2018, Pham.Messai.ea2020a, Pham.Messai.ea2020b, Luo.Ye2022, Zegers.Phillips.ea2021, Nino.Zegers.ea2023] delve into the control of MASs over clustered networks. In [Qin.Yu2013], the authors develop a distributed controller enabling $p$-cluster consensus for a MAS of identical linear time-invariant (LTI) systems coordinating over a static or switching graph. Given a graph encoding a MAS whose vertex set is partitioned into $p$ elements (i.e., clusters), $p$-cluster consensus is said to occur whenever the agents in each cluster achieve consensus, and the consensus value between clusters is distinct. Note that $p$-cluster consensus is accomplished using continuous state feedback. In [Ge.Han.ea2018], the authors investigate the formation control problem for clustered networks of identical LTI systems. The authors develop a distributed controller that enables $M$ clusters to assemble distinct rigid formations using intermittent state feedback while compensating for communication delays.
@@
The result in [Pham.Messai.ea2020a] presents a distributed consensus controller for a clustered network of identical LTI systems, where each agent leverages a mixture of continuous and intermittent output feedback. Each cluster is assigned a single leader, and, within each cluster, all agents may continuously communicate output information with their neighbors. However, only the leaders of each cluster may communicate intermittently with neighboring leaders, and this is the means by which information is spread across clusters. In addition, all leaders communicate with their neighboring leaders at the same isolated times. A Luenberger-like observer produces state estimates from output information, which are then used for control. Albeit not written, the analysis in [Pham.Messai.ea2020a] renders the agreement subspace for the entire ensemble globally exponentially stable (GES). Leveraging analogous arguments and cluster structures found in [Pham.Messai.ea2020a], the result in [Pham.Messai.ea2020b] develops a distributed controller enabling a clustered network of identical LTI systems to assemble a single rigid formation using a mixture of continuous and intermittent state feedback. In [Luo.Ye2022], the authors build on the result of [Qin.Yu2013] and identify the general graph property that supports $p$-cluster consensus independent of coupling strength.
1.3 Comparison to Previous Work
@@
In [Zegers.Phillips.ea2021], we build on [Pham.Messai.ea2020a] and develop a distributed controller for a clustered network of identical LTI systems that enables consensus. Like [Pham.Messai.ea2020a], communication within clusters is continuous while communication between clusters is intermittent. Yet, unlike [Pham.Messai.ea2020a], the result in [Zegers.Phillips.ea2021] does not employ a leader-follower strategy and more general partitions of the communication graph are considered. As a result, more than one agent in a cluster can have a neighbor in a different cluster. Two pairs of distinct clusters may exchange information at different rates (i.e., asynchronously) in [Zegers.Phillips.ea2021], which is not supported in [Pham.Messai.ea2020a]. Still, it is important to note that [Pham.Messai.ea2020a] considers output feedback while [Zegers.Phillips.ea2021] only considers state feedback.
@@
Our result in [Nino.Zegers.ea2023] is an extension of [Zegers.Phillips.ea2021], which develops a distributed controller for a clustered network of identical LTI systems facilitating consensus using only output feedback. Similar to the results in [Pham.Messai.ea2020a] and [Zegers.Phillips.ea2021], communication within clusters is continuous while communication between clusters is intermittent. Rather than broadcast state information between neighbors as in [Zegers.Phillips.ea2021], the result in [Nino.Zegers.ea2023] broadcasts state estimates computed using a Luenberger-like observer instead. This observer leverages intermittently sampled outputs, model knowledge, and control input measurements to compute state estimates.
@@
Unfortunately, the stability analysis in [Nino.Zegers.ea2023] suffers from an artificial defect. To demonstrate the desired set $\mathcal{A}$ is GES, the stability analysis required all Luenberger-like observers to undergo synchronous jumps. The price to pay for this mathematical convenience is that all agents must synchronously sample their output, which is impractical for large-scale networks.
1.4 Control Strategy Description & Contributions
Motivated by our previous work in [Zegers.Phillips.ea2021, Nino.Zegers.ea2023], we introduce a new distributed consensus controller for a clustered network of identical LTI systems that leverages intermittent and asynchronous output feedback both within and between clusters. The clustered network is modeled as a connected, undirected, and static graph, which is decomposed into cluster and inter-cluster sub-graphs. [^1] Such a structure can be used to encode a team of teams, where clusters can describe teams of agents and inter-clusters can describe connections between these teams. Each agent and inter-cluster sub-graph is assigned a unique timer mechanism that dictates when measurement and control update events occur. Each timer mechanism is responsible for generating isolated events times, where the difference between consecutive event times can be bounded above and below by user-defined parameters, allowing for the a priori exclusion of Zeno behavior, the satisfaction of hardware constraints (e.g., sensor sampling rate), and the satisfaction of the maximum allowable transmission interval (MATI) [Phillips.Sanfelice2019]. Within each cluster, each agent intermittently measures the outputs of their same-cluster neighbors and themselves when instructed by their timer mechanism. Similarly, within an inter-cluster, each agent intermittently measures the outputs of their same-inter-cluster neighbors and themselves when instructed by the inter-cluster timer mechanism. Thus, the measurement of output information is completely intermittent throughout the entire network, unlike [Pham.Messai.ea2020a, Zegers.Phillips.ea2021, Nino.Zegers.ea2023]. Since these timer mechanisms are independent, the flow of output information within clusters and inter-clusters may be asynchronous, resulting in the directed flow of output information along the edges of an undirected graph. Every agent computes a consensus-like estimate for their assigned cluster and each inter-cluster the agent is a member of, which evolve in an open-loop fashion during flows and are impulsively reset with the appropriate output measurements during jumps as determined by the corresponding timer mechanism. Each agent uses their consensus-like estimates to compute a control input at each moment in time, which is shown to enable consensus while only using intermittently and asynchronously available output measurements. Such estimators provide a simpler alternative to the Luenberger-like observes utilized in [Nino.Zegers.ea2023], supplanting model-based state reconstruction with model-less output-based estimation.
When considering consensus or synchronization problems over finite-dimensional metric spaces, it is customary to examine disagreement dynamics, where disagreement is usually defined through a projection onto the orthogonal complement of the agreement subspace. However, such a projection is not full rank and can prevent the satisfaction of sufficient conditions expressed as matrix inequalities ensuring the stability of the agreement subspace. To overcome this challenge, we develop an alternative disagreement metric based on the image of a full-rank operator. This alternative disagreement metric is used to derive a closed-loop disagreement dynamics, consisting of continuous-time and discrete-time behaviors, modeled as a hybrid system of the form in [Goebel.Sanfelice.ea2012]. The consensus objective is cast as a set stabilization problem, where the proposed hybrid system is shown to be nominally well-posed, the desired set is shown to be compact and GES for the proposed hybrid system via a Lyapunov-based stability analysis, and the desired set is shown to be robust to Lebesgue measurable and locally essentially bounded perturbations through an input-to-state stability argument. [^2] The nominal well-posedness of the proposed hybrid system, compactness of the desired set, and the GES of the desired set imply this set is robust to vanishing perturbations. Such results are not considered in [Pham.Messai.ea2020a, Zegers.Phillips.ea2021, Nino.Zegers.ea2023] nor is the GES of a set encoding the consensus objective achieved by a control strategy leveraging only intermittent and asynchronous information. To the best of our knowledge, the contributions of this work to the literature are the following:
  1. 1.
    A distributed consensus controller for clustered networks of identical LTI systems utilizing only intermittent and asynchronous output feedback;
  2. 2.
    Stemming from Item 1, a nominally well-posed hybrid system with complete non-Zeno solutions and a GES compact attractor;
  3. 3.
    Certificates of robustness to vanishing perturbations via standard results in [Goebel.Sanfelice.ea2012] and Lebesgue measurable and locally essentially bounded perturbations via an input-to-state stability proof.
The theoretical development is validated in simulation, where we provide an online path planner for a rendezvous application.
1.5 Organization
@@
Section 2 provides information on notation, graphs, cluster and inter-cluster sub-graphs, and hybrid systems. Section 3 describes the specific problem we are aiming to solve. The proposed solution and closed-loop hybrid system are presented in Section 4. In Section 5, we provide the main results of the paper in the form of sufficient conditions for the nominal system and, in Section 6, we extend these results to consider a class of perturbations. Lastly, we illustrate the results through numerical simulations in Section 7.
2 Preliminaries
2.1 Notation
@@
Let $\mathbb{R}$ and $\mathbb{Z}$ denote the set of reals and integers, respectively. For a constant $a\in\mathbb{R}$, let $\mathbb{R}_{\geq a}\coloneqq[a,\infty)$, $\mathbb{R}_{>a}\coloneqq(a,\infty)$, $\mathbb{Z}_{\geq a}\coloneqq\mathbb{R}_{\geq a}\cap\mathbb{Z}$, and $\mathbb{Z}_{>a}\coloneqq\mathbb{R}_{>a}\cap\mathbb{Z}$. For $p,q\in\mathbb{Z}_{>0}$, the $p\times q$ zero matrix and the $p\times 1$ zero column vector are respectively denoted by $0_{p\times q}$ and $0_{p}$. The $p\times p$ identity matrix and the $p\times 1$ column vector with all entries being one are denoted by $I_{p}$ and $1_{p}$, respectively. The Euclidean norm of $r\in\mathbb{R}^{p}$ is denoted by $\|r\|\coloneqq\sqrt{r^{\top}r}$. For $x,y\in\mathbb{R}^{p}$, the inner product between $x$ and $y$ is denoted by $\langle x,y\rangle\coloneqq x^{\top}y$. For $M\in\mathbb{Z}_{\geq 2}$, let $M]\coloneqq\{1,2,...,M\}$. The Kronecker product between $A\in\mathbb{R}^{p\times q}$ and $B\in\mathbb{R}^{u\times v}$, where $u,v\in\mathbb{Z}_{>0}$, is denoted by $A\otimes B\in\mathbb{R}^{pu\times qv}$. The maximum and minimum eigenvalues of a real symmetric matrix $A\in\mathbb{R}^{n\times n}$ are denoted by $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$, respectively. The block diagonal matrix with general blocks $G_{1},G_{2},...,G_{p}$ is denoted by $\text{diag}(G_{1},G_{2},...,G_{p})$. The distance of a point $r\in\mathbb{R}^{p}$ to the set $S\subset\mathbb{R}^{p}$ is given by $|r|_{S}\coloneqq\inf\{\|r-s\|\colon s\in S\}\in\mathbb{R}_{\geq 0}$, and let $r+S\coloneqq\{r+s\in\mathbb{R}^{p}\colon s\in S\}$. For a collection of vectors $\{z_{1},z_{2},...,z_{p}\}\subset\mathbb{R}^{q}$, let $(z_{k})_{k\in[p]}\coloneqq[z_{1}^{\top},z_{2}^{\top},...,z_{p}^{\top}]^{\top}\in\mathbb{R}^{pq}$. Similarly, for $x\in\mathbb{R}^{p}$ and $y\in\mathbb{R}^{q}$, let $(x,y)\coloneqq[x^{\top},y^{\top}]^{\top}\in\mathbb{R}^{p+q}$. For any sets $A$ and $B$, a function $f$ of $A$ with values in $B$ is denoted by $f\colon A\to B$, whereas $f\colon A\rightrightarrows B$ refers to a set-valued function $f\colon A\to 2^{B}$. The $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N}$, for $N\in\mathbb{Z}_{>0}$, is denoted by $\mathtt{e}_{k}$. Similarly, let $\tilde{\mathtt{e}}_{k}$ denote the $k^{\text{th}}$ standard basis vector of $\mathbb{R}^{N-1}$. The range (i.e., image) of a matrix $A\in\mathbb{R}^{N\times N}$ is denoted by $\text{Range}(A)\coloneqq\{y\in\mathbb{R}^{N}\colon y=Ax\text{ for some }x\in\mathbb{R}^{N}\}$. For any $x\in\mathbb{R}$, let $\lfloor x\rfloor\coloneqq\sup\{c\in\mathbb{Z}\colon c\leq x\}$. Given a symmetric real-valued matrix $A$, the notation $A>\mathbf{0}$, $A\geq\mathbf{0}$, $A<\mathbf{0}$, and $A\leq\mathbf{0}$ imply $A$ is positive definite, positive semi definite, negative definite, and negative semi definite, respectively.
2.2 Graphs